OpenStudy (anonymous):

What polynomial has roots of -4, -1, and 6 ? x3 - 9x2 - 22x + 24 x3 - x2 - 26x - 24 x3 + x2 - 26x + 24 x3 + 9x2 + 14x - 24

mathslover (mathslover):

The best way to do this question will be Hit and Trial Method... Put the values (-4,-1,-6) at the place of x in all the expressions, the one which equates to zero is your answer.

hero (hero):

If a polynomial has roots of -4, -1, and 6, Then: x = -4 x = -1 x = 6 Thus: x + 4 = 0 x + 1 = 0 x - 6 = 0 By zero product property: (x + 4)(x - 1)(x - 6) = 0 Expand (x + 4)(x - 1)(x - 6) to find the correct polynomial